# Planck dust polarization power spectra are consistent with strongly supersonic turbulence

### v1

@@ Abstract
The polarization of the Cosmic Microwave Background (CMB) is rich in information but obscured by foreground emission from the Milky Way’s interstellar medium (ISM). To uncover relationships between the underlying turbulent ISM and the foreground power spectra, we simulated a suite of driven, magnetized, turbulent models of the ISM, varying the fluid properties via the sonic Mach number, $\mathcal{M_{S}}$, and magnetic (Alfvén) Mach number, $\mathcal{M_{\rm{A}}}$. We measure the power spectra of density ($\rho$), velocity ($v$), magnetic field ($H$), total projected intensity ($T$), parity-even polarization ($E$), and parity-odd polarization ($B$). We find that the slopes of all six quantities increase with $\mathcal{M_{S}}$ and tend to increase with $\mathcal{M_{\rm{A}}}$. By comparing spectral slopes of $E$ and $B$ to those measured by Planck, we infer typical values of $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ for the ISM. As the fluid velocity increases, the ratio of BB power to EE power increases to approach a constant value near the Planck-observed value of $\sim 0.5$ for $\mathcal{M_{S}}>4$, regardless of the magnetic field strength. We also examine correlation-coefficients between projected quantities, and find that $r^{\rm{TE}}\approx 0.3$, in agreement with Planck, for appropriate combinations of $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$. Finally, we consider parity-violating correlations $r^{\rm{TB}}$ and $r^{\rm{EB}}$.
@@ 1 Introduction
Primordial gravitational waves, generated in inflation and imprinted on the surface of last scattering, are one of the most exciting sources of polarization in the Cosmic Microwave Background (CMB). Their discovery would give revolutionary evidence for inflation and its mechanism (Kamionkowski & Kovetz, 2016). However, the brightest diffuse sources of polarization in the microwave sky are thermal dust and synchrotron emission. These are also exciting signals, because they reveal the turbulent magnetic field in the interstellar medium (ISM) of our own Galaxy (Kritsuk et al., 2018; Kim et al., 2019). In order to see the potential inflation signal, we must first characterize and mitigate the Galactic signal which is at least ten times larger (Planck Collaboration et al., 2020).
@@
Linearly polarized light can be described by the Stokes $Q,U$ parameters, which characterize the propagation of polarized light, but these quantities are coordinate dependent. The coordinate-independent $E,B$ parameters, on the other hand, convert the $Q,U$ signal into parity-even $E$ modes and parity-odd $B$ modes. At the surface of last scattering, cosmological scalar density perturbations produce $E$, while in the filamentary ISM, the $E$ modes are strongest for polarization parallel or perpendicular to a filament or sheet. On the other hand, at the surface of last scattering, gravitational waves (cosmological tensor perturbations) are the only producer of primordial $B$-mode polarization, peaking at degree scales and larger, while in the ISM, the $B$-modes are strongest for polarization at an oblique angle to a filament (Huffenberger et al., 2020). As the CMB signal travels from the surface of last scattering along the line of sight, gravitational lensing by large scale structure also generates $B$-modes, and CMB-lensing reconstruction could also be contaminated by ISM signals.
@@
The Planck satellite measured $E$-mode and $B$-mode power spectra at 353 GHz, where the signal is dominated by dust emission, and found approximate powerlaws with slopes of $\alpha_{EE}=-2.42\pm 0.02$ and $\alpha_{BB}=-2.54\pm 0.02$ for a 71% sky area (Planck Collaboration et al., 2020). For the same sky area, the ratio of amplitudes of $B$ power to $E$ power was $0.53\pm 0.01$. They also found that the scalar temperature $T$-mode and $E$-mode are correlated with correlation coefficient $r^{TE}=0.36$ with some scatter but no clear trend with sky region or with angular scale. Curiously, the $T$-modes and $B$-modes are correlated with $r^{TB}=0.05$. This was unexpected as the correlation of a parity-even $T$-mode and a parity-odd $B$-mode should be zero on average in systems with no helicity or parity violation, and may indicate something about the structure of the magnetic field in the solar neighborhood (Planck Collaboration et al., 2020). Huffenberger et al. (2020) pointed out that in a filamentary picture, a positive $TB$ dust correlation would imply a positive $EB$ correlation, one that is too small to detect with Planck data alone, but is amenable to searches assisted with ISM tracers like HI (Cukierman et al., 2023).
@@
In addition to searches for inflation, the polarization of the CMB can be used to hunt for signatures of other beyond-the-standard-model physics. Cosmic birefringence is the rotation of the polarization vector of CMB photons by hypothetical pseudoscalar fields. Possibilities include axion-like particles that may be responsible for dark matter (Komatsu, 2022). Minami & Komatsu (2020a) report the measurement of a rotation angle of $\beta=(0.35\pm 0.14)^{\circ}$ due to cosmic birefringence. In their work, they assume that the $E$ and $B$ polarization produced by the ISM are uncorrelated in order to set the calibration for the polarization angle. Because the sign of their correlation is the opposite of what is implied for dust in the filament picture, a nonzero $EB$ correlation in the dust would make their result stronger. (Minami & Komatsu, 2020b; Diego-Palazuelos et al., 2022, 2023, provide refinements and robustness tests.)
@@
We organize our paper as follows. We describe the methods for simulations and analysis in Section 2. We present results, beginning with a short overview, in Section 3. There we show power spectra and slopes for both fluid (Section 3.1) and projected quantities (Section 3.2). We present ratios and correlation coefficients (Section 3.3). Using the measured quantities from Planck, we posit typical values for the ISM’s sonic and Alfvén Mach numbers (Section 3.5). We briefly contrast these findings with projections parallel to the mean magnetic field (Section 3.6). We conclude in Section 4.
Refer to captionRefer to captionRefer to caption
@@ 2.1 Simulations
The simulation boxes are periodic, use a $512^{3}$-zone resolution, and start with uniform density. To generate turbulence, we drive the fluid with a stochastic forcing that adds a random acceleration pattern
[…]
@@
where $\rho$ is the mean density, $v_{A}$ is the Alfvén velocity, $c_{s}$ is the speed of sound, and $H$ is the mean magnetic field strength. Stronger field means bigger $v_{A}$ and thus smaller $\mathcal{M_{\rm{A}}}$.
Throughout the work we define $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ with the 3D velocity dispersion, $v_{\rm rms}^{2}=\langle v_{x}^{2}\rangle+\langle v_{y}^{2}\rangle+\langle v_{z}^{2}\rangle$. In contrast to the 3D Mach number $\mathcal{M_{S}}$, the 1D Mach number (which enters the Maxwellian velocity distribution (Rabatin & Collins, 2023) and is more accessible from the ground) is found by assuming isotropy and dividing $v_{\rm rms}$ by $\sqrt{3}$.
@@
Figure 1 shows $\ln(T)$, $E$, and $B$ (defined precisely in the following section). Each figure shows three target Sonic Mach numbers, (0.5, 3, 6) and Alfvén Mach numbers (0.5, 1, 2). The colorbar is a symmetric logarithm. Clearly, structure changes as $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ increase. This is due to the fact that the dissipation scale decreases as the mean kinetic energy increases.
Figure 2 shows a legend of the achieved $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ for each simulation. Because it is challenging to predict the forcing to achieve particular Mach numbers, the measured values of $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ differ slightly from the nominal values listed above. This figure also serves as a legend for the remaining figures in the work, with color denoting mach number, with dark blue-to-yellow corresponding to achieved $\mathcal{M_{S}}$, ranging from subsonic ($\mathcal{M_{S}}<1$) to supersonic ($\mathcal{M_{S}}\sim 7$).
@@ 2.3 Power spectra
where $\tilde{X}$ and $\tilde{Y}$ are Fourier transforms of fluid quantities ($\rho$, $\mathbf{v}$, and $\mathbf{H}$, whence dimension $D=3$) or projected quantities ($T$, $E$, and $B$, whence $D=2$). $\Delta V_{k}$ is the volume of a shell at $k$, which has thickness matched to the resolution of the Fourier grid, $\Delta k=k_{\rm min}=2\pi/L$. Because the box is scale-free, the wavenumbers $k$ do not correspond to any particular angular scale or multipole.
@@ 3 Results
Figure 3 gives an overview of the main result for foreground polarization. The vertical axis shows $\alpha_{EE}$ while the horizontal shows $A_{\rm{BB}}/A_{\rm{EE}}$ (the ratio of the fit amplitudes). Grey lines indicate the Planck-measured values. The color shows the Mach number $\mathcal{M_{S}}$ and the size shows $\mathcal{M_{\rm{A}}}$ (as in Fig. 2). As discussed below, as $\mathcal{M_{S}}$ increases above 4, the ratio $A_{\rm{BB}}/A_{\rm{EE}}$ increases to sit in the range [0.4,0.7], compared to the Planck value near 0.5, and $\alpha_{EE}$ becomes shallower to a range $[-2.6,-2.2]$, compared to the Planck value of -2.42.
Figure 4: Power spectra of density ($C_{k}^{\rho\rho}$), velocity ($C_{k}^{vv}$) and magnetic field strength ($C_{k}^{HH}$) (left to right, respectively.) All plots have been compensated to the Kolmogorov value of 11/3. Color (blue-yellow) denotes increasing Mach number, while linestyle denotes $\mathcal{M_{\rm{A}}}$ as described in Section 2.1: dotted lines for $0.70.4<\mathcal{M_{\rm{A}}}<$ 0.7, dashed lines for $1.20.7<\mathcal{M_{\rm{A}}}<$ 1.2, and solid lines for $2.21.2<\mathcal{M_{\rm{A}}}<$ 2.2. Density slope depends heavily on Mach number, while velocity and magnetic slopes do not.
Figure 5: Slopes of fluid quantities ($\rho$, ${\bf v}$, and ${\mathbf{H}}$). The first column of each plot shows (slopes, amplitdues) relative to $\mathcal{M_{S}}$ for the first two, and $\mathcal{M_{\rm{A}}}$ for $\alpha_{HH}$. Color denotes $\mathcal{M_{S}}$ while size denotes $\mathcal{M_{\rm{A}}}$ as described in Section 2.1.
@@ 3.1 Fluid Power Spectra
Figures 4 display the spectra for the fluid quantities, $C_{k}^{\rho\rho}$ (density, left), $C_{k}^{{vv}}$(velocity, center) and $C_{k}^{HH}$ (magnetic field, right). Figure 5 shows the slopes of those spectra, $\alpha_{\rho\rho}$ (left), $\alpha_{{vv}}$ (center) and $\alpha_{HH}$ (right). The spectra are compensated by $k^{11/3}$ to emphasize variations relative to the average Kolmogorov slope value, which would be flat in this plot. The plot style is described in Section 2.1; color (blue to yellow) denotes increasing $\mathcal{M_{S}}$, while line style denotes $\mathcal{M_{\rm{A}}}$. Figure 5 shows the slopes, $\alpha_{XX}$, with color denoting $\mathcal{M_{S}}$ and point size increasing with $\mathcal{M_{\rm{A}}}$.
@@
Beginning with density (first panels in Figures 4 and 5) we find that both slope and amplitude are increasing functions of $\mathcal{M_{S}}$. The increase in amplitude is to be expected, as the variance in density is linearly proportional to the variance in velocity (Federrath et al., 2008), i.e.
$\sigma^{2}_{\rho}=b^{2}\sigma^{2}_{v},$ (13)
where $b\in[1/3,1]$. By the Plancherel theorem $\sigma^{2}_{\rho}$ is an integral of the power spectrum, so it is expected that the amplitude of the density power spectra will increase with Mach number. The slope becomes shallower in an almost linear way from $\alpha_{\rho\rho}\sim-3.5$ to $\alpha_{\rho\rho}\sim-2.5$ as $\mathcal{M_{S}}$ increases from 0.5 to 7. As $\mathcal{M_{S}}$ increases, the typical shock velocity also increases, which gives rise to enhanced structure formation by way of fluid instabilities such as the Richtmyer-Meshkov instability (Richtmyer, 1960; Meshkov, 1972). This enhanced power flattens the spectrum. This behavior has been seen before (Beresnyak et al., 2005; Collins et al., 2012). Neither slope, $\alpha_{\rho\rho}$ nor amplitude, $A_{\rho}$ vary with Alfvén Mach number. This is not particularly surprising, as the continuity equation which determines density only contains density and velocity:
$\frac{\partial\rho}{\partial t}+\nabla\cdot{\bf v}\rho=0,$
so the density is determined by velocity. We see that the density power spectrum is apparently determined solely by way of the r.m.s. velocity.
@@
The velocity field (second panels of Figures 4 and 5) can be seen to vary jointly with $\mathcal{M_{\rm{A}}}$ and $\mathcal{M_{S}}$. In incompressible hydrodynamical turbulence, the expectation is that the slope has the (average) Kolmogorov value of $-11/3\approx-3.7$. These simulations are not imcompressible, but highly compressible and magnetized. For supersonic hydrodynamical turbulence, one expects a value of $\alpha_{{vv}}=-4$ or ($\alpha_{\rm{total}}=-2$, Kritsuk et al. 2007.) For incompressible magnetized turbulence, we expect a value of $\alpha_{v}=-5/2$ along the field, and $\alpha_{v}=-11/3$ transverse to the magnetic field ($\alpha_{\rm{total}}=-3/2$ and $-5/3$, respectively, Goldreich & Sridhar 1995). There is not a theory that combines compressibility and magnetization that is appropriate for the simulations presented here, and in our case we see some resemblance to all of the above. The simulation that most closely approaches the un-magnetized incompressible assumption of the Kolmogorov cascade has $\mathcal{M_{S}}$$=0.5$ and $\mathcal{M_{\rm{A}}}$$=2$, which does have a slope of $-3.6$. For low $\mathcal{M_{S}}$, the slope $\alpha_{{vv}}$ steepens from $-3.6$ to $-3.9$ as the field increases (dot size shrinks). Once supersonic, the slope of the velocity does not vary much with $\mathcal{M_{S}}$, but does steepen with increasing field strength. For low field strength, the slope is around around $\alpha_{{vv}}=-3.5$ ($\alpha_{\rm{total}}=-1.5$).
@@
The final fluid quantity is magnetic field, ${\mathbf{H}}$. Spectra are plotted in Figure 4, and slopes are plotted in Figure 5). Here the magnetic slope, $\alpha_{HH}$ is plotted against Alfvén Mach number rather than Sonic Mach number. It can be seen that the slope of the magnetic field, $\alpha_{HH}$, does not depend strongly on $\mathcal{M_{S}}$, as points with similar color cluster around the same value, but does decrease nearly linearly for decreasing magnetic field strength. For the weakly magnetized runs, $\alpha_{HH}=-3.75$, and it increases to $-3.3$ for the strongly magnetized runs. This is an intuitively reasonable result, as the magnetic field strength increases ($\mathcal{M_{\rm{A}}}$ decreases) the ability of the field to resist tangling by kinetic motions increases, which suppresses power at all scales, but especially large scales, making the slope more shallow.

### v2

@@ Abstract
The polarization of the Cosmic Microwave Background (CMB) is rich in information but obscured by foreground emission from the Milky Way’s interstellar medium (ISM). To uncover relationships between the underlying turbulent ISM and the foreground power spectra, we simulated a suite of driven, magnetized, turbulent models of the ISM, varying the fluid properties via the sonic Mach number, $\mathcal{M_{S}}$, and magnetic (Alfvén) Mach number, $\mathcal{M_{\rm{A}}}$. We measure the power spectra of density ($\rho$), velocity ($v$), magnetic field ($H$), total projected intensity ($T$), parity-even polarization ($E$), and parity-odd polarization ($B$). We find that the slopes of all six quantities increase with $\mathcal{M_{S}}$. Most increase with $\mathcal{M_{\rm{A}}}$, while the magnetic field spectrum steepens with $\mathcal{M_{\rm{A}}}$. By comparing spectral slopes of $E$ and $B$ to those measured by Planck, we infer typical values of $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ for the ISM. As the fluid velocity increases, $\mathcal{M_{S}}>4$, the ratio of BB power to EE power increases to approach a constant value near the Planck-observed value of $\sim 0.5$, regardless of the magnetic field strength. We also examine correlation-coefficients between projected quantities, and find that $r^{\rm{TE}}\approx 0.3$, in agreement with Planck, for appropriate combinations of $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$. Finally, we consider parity-violating correlations $r^{\rm{TB}}$ and $r^{\rm{EB}}$.
@@ 1 Introduction
Primordial gravitational waves, generated during inflation and imprinted on the surface of last scattering, are one of the most exciting sources of polarization in the Cosmic Microwave Background (CMB). Their discovery would give revolutionary evidence for inflation and its mechanism (Kamionkowski & Kovetz, 2016). However, the brightest diffuse sources of polarization in the microwave sky are thermal dust and synchrotron emission. These are also exciting signals, because they reveal the turbulent magnetic field in the interstellar medium (ISM) of our own Galaxy (Planck Collaboration et al., 2015; Kritsuk et al., 2018; Kim et al., 2019). In order to see the potential inflation signal, we must first characterize and mitigate the Galactic signal which is at least ten times larger (Planck Collaboration et al., 2020).
@@
Linearly polarized light can be described by the Stokes $Q,U$ parameters, but these quantities are coordinate dependent. The coordinate-independent $E,B$ parameters, on the other hand, transform the $Q,U$ signal into parity-even $E$ modes and parity-odd $B$ modes. The transformation is nonlocal, and size of the kernel depends on the band limit (Rotti & Huffenberger, 2019). At the surface of last scattering, cosmological scalar density perturbations produce $E$, while gravitational waves (cosmological tensor perturbations) are the only producer of primordial $B$-mode polarization, peaking at degree scales and larger. As the CMB photons travels from the surface of last scattering along the line of sight, gravitational lensing by large scale structure also generates $B$-modes from the scalar $E$ modes. All of these encode rich information about the physics and cosmology of the Universe.
In the ISM, $E$ and $B$ (because they are nonlocal) depend on the geometry of structures and their spatial relationship to the magnetic field (which sets the polarization directions). Gravitational, pressure, and magnetic forces can produce ISM structures with one long and two short dimensions, so-called filaments, and such filamentary structures are apparent in millimeter and HI data (Clark et al., 2014; Planck Collaboration et al., 2016; Clark & Hensley, 2019) as well as ISM simulations (de Avillez & Breitschwerdt, 2005; Hennebelle, 2013). A filament will produce predominantly $E$ modes when paired with a magnetic field parallel or perpendicular to the long axis, and a filament will produce predominantly $B$ modes when paired with an oblique magnetic field (Huffenberger et al., 2020). We may thus expect that a sufficiently strong magnetic field (that aligns filaments to its direction) will produce less $B$ signal than $E$.
@@
The Planck satellite measured $E$-mode and $B$-mode power spectra at 353 GHz, where the signal is dominated by dust emission in the ISM, and found approximate powerlaws with slopes of $\alpha_{EE}=-2.42\pm 0.02$ and $\alpha_{BB}=-2.54\pm 0.02$ for a 71% sky area (Planck Collaboration et al., 2020). For the same sky area, the ratio of amplitudes of $B$ power to $E$ power was $0.53\pm 0.01$. They also found that the scalar temperature $T$-mode and $E$-mode are correlated with correlation coefficient $r^{TE}=0.36$, with some scatter but no clear trend with sky region or with angular scale. Curiously, the $T$-modes and $B$-modes are correlated with $r^{TB}=0.05$. This was unexpected as the correlation of a parity-even $T$-mode and a parity-odd $B$-mode should be zero on average in systems with no helicity or parity violation, and may indicate something about the structure of the magnetic field in the solar neighborhood (Planck Collaboration et al., 2020). Huffenberger et al. (2020) pointed out that in a filamentary picture, a positive $TB$ dust correlation would imply a positive $EB$ correlation. Such a signature is too small to detect with Planck data alone, but is amenable to searches assisted with ISM tracers like HI (Cukierman et al., 2023).
@@
In addition to searches for inflation, the polarization of the CMB can be used to hunt for signatures of other beyond-the-standard-model physics. Cosmic birefringence is the rotation of the polarization of CMB photons by hypothetical pseudoscalar fields. Possibilities include axion-like particles that may be responsible for dark matter (Komatsu, 2022). Minami & Komatsu (2020a) employed a strategy to measure the CMB’s $EB$ correlation, using the foreground $EB$ correlation to calibrate detector polarization angles. Based on the CMB $EB$ correlation, they report a rotation angle of $\beta=(0.35\pm 0.14)^{\circ}$ due to cosmic birefringence, when they assume that the foreground $E$ and $B$ polarization produced by the ISM are uncorrelated. If instead the sign of the foreground $EB$ is assumed positive (as implied in the filament picture above or for any other reason), their result indeed gets stronger. (Minami & Komatsu, 2020b; Diego-Palazuelos et al., 2022, 2023, provide refinements and robustness tests to this approach.)
@@
We organize our paper as follows. We describe the methods for simulations and analysis in Section 2. We present results, beginning with a short overview, in Section 3. There we show power spectra and slopes for both fluid (Section 3.1) and projected quantities (Section 3.2). We present ratios of $T$, $E$, and $B$ and thier correlation coefficients (Section 3.3). Using the measured quantities from Planck, we posit typical values for the ISM’s sonic and Alfvén Mach numbers (Section 3.5). We briefly contrast these findings with projections parallel to the mean magnetic field (Section 3.6). We discuss the relation between these findings and our filamentary model (Section 4). We conclude in Section 5.
@@ 2.1 Simulations
where $\rho$ is the mean density, $v_{A}$ is the Alfvén velocity, $c_{s}$ is the speed of sound, and $H$ is the mean magnetic field strength. Stronger field means bigger $v_{A}$ and thus smaller $\mathcal{M_{\rm{A}}}$. In these simulations we drive turbulence in periodic boxes, altering the sonic and Alfvén Mach numbers, and explore the resulting power spectra and correlations.
[…]
In practice, the mean density and sound speed are both set to unity, and the Mach number is controlled by setting the energy injection rate, $\dot{E}\propto\mathcal{M_{S}}^{3}/L$.
Throughout the work we define $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ with the 3D velocity dispersion, $v_{\rm rms}^{2}=\langle v_{x}^{2}\rangle+\langle v_{y}^{2}\rangle+\langle v_{z}^{2}\rangle$. In contrast to the 3D Mach numbers, the 1D Mach number (which enters the Maxwellian velocity distribution (Rabatin & Collins, 2023) and is more accessible from the ground) is found by assuming isotropy and dividing $v_{\rm rms}$ by $\sqrt{3}$.
@@
Figure 1 shows $\ln(T)$, $E$, and $B$ (defined precisely in the following section). Each figure shows three target sonic Mach numbers, (0.5, 3, 6) and Alfvén Mach numbers (0.5, 1, 2). The colorbar is a symmetric logarithm. Structure changes can clearly be seen as $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ increase. This is due to the fact that the dissipation scale decreases as the mean kinetic energy increases.
Figure 2 shows a legend of the achieved $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ for each simulation. Because it is challenging to predict the forcing to achieve particular Mach numbers, the measured values of $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ differ slightly from the nominal values listed above. This figure also serves as a legend for the remaining figures in the work, with color denoting sonic Mach number, with dark blue-to-yellow corresponding to achieved $\mathcal{M_{S}}$, ranging from subsonic ($\mathcal{M_{S}}<1$) to supersonic ($\mathcal{M_{S}}\sim 7$).
@@ 2.3 Power spectra
where $\tilde{X}$ and $\tilde{Y}$ are Fourier transforms of fluid quantities ($\rho$, $\mathbf{v}$, and $\mathbf{H}$, whence dimension $D=3$) or projected quantities ($T$, $E$, and $B$, whence $D=2$). $\Delta V_{k}$ is the volume of a shell at $k$, which has thickness matched to the resolution of the Fourier grid, $\Delta k=k_{\rm min}=2\pi/L$. Because the box is scale-free, the wavenumbers $k$ do not correspond to any particular angular scale or multipole on the sky.
@@ 3 Results
Figure 3 gives an overview of the main result for foreground polarization. The vertical axis shows $\alpha_{EE}$ while the horizontal shows $A_{\rm{BB}}/A_{\rm{EE}}$ (the ratio of the fit amplitudes). Grey lines indicate the Planck-measured values. The color shows the sonic Mach number $\mathcal{M_{S}}$ and the size shows $\mathcal{M_{\rm{A}}}$ (as in Fig. 2). As discussed below, as $\mathcal{M_{S}}$ increases above 4, the ratio $A_{\rm{BB}}/A_{\rm{EE}}$ increases to sit in the range [0.4,0.7], compared to the Planck value near 0.5, and $\alpha_{EE}$ becomes shallower to a range $[-2.6,-2.2]$, compared to the Planck value of -2.42.
Figure 4: Power spectra of density ($C_{k}^{\rho\rho}$), velocity ($C_{k}^{vv}$) and magnetic field strength ($C_{k}^{HH}$) (left to right, respectively.) All plots have been compensated to the Kolmogorov value of 11/3. Color (blue-yellow) denotes increasing sonic Mach number, while line style denotes $\mathcal{M_{\rm{A}}}$ as described in Section 2.1: dotted lines for $0.70.4<\mathcal{M_{\rm{A}}}<$ 0.7, dashed lines for $1.20.7<\mathcal{M_{\rm{A}}}<$ 1.2, and solid lines for $2.21.2<\mathcal{M_{\rm{A}}}<$ 2.2. Density slope depends heavily on sonic Mach number, while velocity and magnetic slopes do not.
Figure 5: Slopes of fluid quantities ($\rho$, ${\bf v}$, and ${\mathbf{H}}$). The first two are relative to $\mathcal{M_{S}}$ and $\mathcal{M_{\rm{A}}}$ is relative to $\alpha_{HH}$. Color denotes $\mathcal{M_{S}}$ while size denotes $\mathcal{M_{\rm{A}}}$ as described in Section 2.1.
@@ 3.1 Fluid Power Spectra
Figures 4 displays the spectra for the fluid quantities, $C_{k}^{\rho\rho}$ (density, left), $C_{k}^{{vv}}$(velocity, center) and $C_{k}^{HH}$ (magnetic field, right). Figure 5 shows the slopes of those spectra, $\alpha_{\rho\rho}$ (left), $\alpha_{{vv}}$ (center) and $\alpha_{HH}$ (right). The spectra are compensated by $k^{11/3}$ to emphasize variations relative to the average Kolmogorov slope value, which would be flat in this plot. The plot style is described in Section 2.1; color (blue to yellow) denotes increasing $\mathcal{M_{S}}$, while line style denotes $\mathcal{M_{\rm{A}}}$. Figure 5 shows the slopes, $\alpha_{XX}$, with color denoting $\mathcal{M_{S}}$ and point size increasing with $\mathcal{M_{\rm{A}}}$. We will discuss each in turn.
@@
Beginning with density (first panels in Figures 4 and 5) we find that both slope and amplitude are increasing functions of $\mathcal{M_{S}}$. The increase in amplitude is to be expected, as the variance in density is linearly proportional to the variance in velocity (Padoan et al., 1997), i.e.
$\sigma^{2}_{\rho}=b^{2}\mathcal{M_{S}}^{2},$ (13)
where $b\in[1/3,1]$ (Federrath et al., 2008). By the Plancherel theorem the variance is equal to the area under the power spectrum, so it is expected that the amplitude of the density power spectra will increase with sonic Mach number. The slope becomes shallower in an almost linear way from $\alpha_{\rho\rho}\sim-3.5$ to $\alpha_{\rho\rho}\sim-2.5$ as $\mathcal{M_{S}}$ increases from 0.5 to 7. As $\mathcal{M_{S}}$ increases, the typical shock velocity also increases, which gives rise to enhanced structure formation by way of fluid instabilities such as the Richtmyer-Meshkov instability (Richtmyer, 1960; Meshkov, 1972). This enhanced power flattens the spectrum. This behavior has been seen before (Beresnyak et al., 2005; Collins et al., 2012).
Neither slope, $\alpha_{\rho\rho}$ nor amplitude, $A_{\rho}$ vary with Alfvén Mach number. This is not particularly surprising, as the continuity equation which determines density only contains density and velocity:
$\frac{\partial\rho}{\partial t}+\nabla\cdot{\bf v}\rho=0,$ (14)
so the density is determined by velocity. Specifically, we find that $\alpha_{\rho\rho}$ is determined by the r.m.s. velocity alone. However, as we will discuss in the next section, the velocity spectrum does not evolve with sonic Mach number in the same manner as the density spectrum, indicating that the shape of the density spectrum is not a direct result of the shape of the velocity spectrum.
@@
The velocity field (second panels of Figures 4 and 5) can be seen to vary jointly with $\mathcal{M_{\rm{A}}}$ and $\mathcal{M_{S}}$. In incompressible hydrodynamical turbulence, the expectation is that the slope has the (average) Kolmogorov value of $-11/3\approx-3.7$. These simulations are not imcompressible, but highly compressible and magnetized. For supersonic hydrodynamical turbulence, one expects a value of $\alpha_{{vv}}=-4$ or ($\alpha_{\rm{total}}=-2$, Kritsuk et al. 2007.) For incompressible magnetized turbulence, the spectral scaling has been debated, with some authors expecting a value of $\alpha_{{vv}}=-5/2$ ($\alpha_{\rm{total}}$=$-3/2$ Iroshnikov, 1964; Boldyrev, 2006)) and some expecting a value of $-11/3$ ($\alpha_{\rm{total}}$=$-5/3$ Beresnyak, 2011). For a recent overview, see Schekochihin (2022). There is not a theory that combines compressibility and magnetization that is appropriate for the simulations presented here, and in our case we see some resemblance to all of the above. The simulation that most closely approaches the un-magnetized incompressible assumption of the Kolmogorov cascade has $\mathcal{M_{S}}$$=0.5$ and $\mathcal{M_{\rm{A}}}$$=2$, which does have a slope of $-3.6$. For low $\mathcal{M_{S}}$, the slope $\alpha_{{vv}}$ steepens from $-3.6$ to $-3.9$ as the field increases (dot size shrinks). Once supersonic, the slope of the velocity does not vary much with $\mathcal{M_{S}}$, but does steepen with increasing field strength. For low field strength, the slope is around around $\alpha_{{vv}}=-3.5$ ($\alpha_{\rm{total}}=-1.5$). Also shown in the figure are horizontal lines showing the fiducial values of $\alpha_{{vv}}=-5/2$ ($\alpha_{\rm{total}}=-3/2$) and $\alpha_{{vv}}=-11/3$ ($\alpha_{\rm{total}}=-5/3$).
@@
The final fluid quantity is magnetic field, ${\mathbf{H}}$. Spectra are plotted in the last panel of Figure 4, and slopes are plotted in the last panel of Figure 5). Here the magnetic slope, $\alpha_{HH}$ is plotted against Alfvén Mach number rather than sonic Mach number. It can be seen that the slope of the magnetic field, $\alpha_{HH}$, does not depend strongly on $\mathcal{M_{S}}$, as points with similar color cluster around the same value, but does decrease nearly linearly for decreasing magnetic field strength. For the weakly magnetized runs, the slope is $\alpha_{HH}=-3.75$, and it becomes shallower to $-3.3$ for the strongly magnetized runs.
This behavior is likely the compressible analog of the transition from weak turbulence, where magnetic fluctuations are smaller than the mean, to strong turbulence, where the fluctuations are large compared to the mean magnetic field (see Schekochihin, 2022, for an excellent review). In incompressible simulations, the spectrum is observed to steepen continually from $\alpha_{\rm{total}}=-3/2$ to $\alpha_{\rm{total}}=-2$ as the turbulence move from weak to strong (Perez & Boldyrev, 2008). This is similar to the monotonic steepening in $\alpha_{HH}$ as we increase $\mathcal{M_{\rm{A}}}$, which seems to level off above $\mathcal{M_{\rm{A}}}\sim 1.5$. In addition to the transition from weak to strong turbulence as $\mathcal{M_{\rm{A}}}$ increases, in our simulations we are also transitioning to shock dominated turbulence, which complicates the picture relative to the incompressible work.